# Nonequilibrium Probes of Quantum Geometry in Gapless Systems

### v1

@@ 1.3 Plan of the paper
The organization of the paper reflects the thematic split outlined above. In Sec. 2, we introduce the family of CFT Hamiltonians (1.1) and their quantum geometry. Sec. 3 relates this to observable results (such as transition rates and linear-response coefficients) for small perturbations (1.3). Sec. 4 concerns large but adiabatically driven deformations, and the general behavior of the return probability exemplified in Fig. 1 is derived for the important class of rotating drives. In Sec. 5, we present families of time-dependent deformations that can be solved exactly using finite-dimensional matrices, without resorting to expansions in different limiting regimes. We conclude in Sec. 6, listing some natural extensions and follow-ups. Some technical details are deferred to appendices.
2 Virasoro quantum geometry
@@
This section introduces the quantum geometry of the emergent conformal symmetry of gapless quantum systems, phrased within the context of CFT. We first recall some basics of conformal transformations and Virasoro group theory, then introduce inhomogeneous CFT and coherent states, and finally turn to our main act: the quantum geometry of CFT.
2.1 Virasoro group theory
@@
We begin by reviewing conformal transformations in any 1+1D CFT. The presentation starts from the group of conformal transformations and introduces their algebra as a second step. [^4] Due to its infinite dimension, representations of the Virasoro group should strictly speaking be built the other way around: starting from the algebra and integrating to the group [54, 55]. We deliberately deviate from this, since our presentation allows one to directly generalize the approach to any quantum system with a (finite-dimensional) continuous symmetry group, where properties such as existence are less of an issue. Select formulas are summarized in Appendix A.
@@
$\varphi^{-}\mapsto f(\varphi^{-}),\qquad\varphi^{+}\mapsto\bar{f}(\varphi^{+}),$ (2.1)
@@
$f^{\prime}(\varphi)>0\qquad\text{and}\qquad f(\varphi+2\pi)=f(\varphi)+2\pi,$ (2.2)
and similarly for $\bar{f}$. The group operation in $\operatorname{Diff}(S^{1})$ is given by composition $(f\circ g)(\varphi)=f(g(\varphi))$, the identity element is $\mathbb{I}(\varphi)=\varphi$, and $f^{-1}$ denotes the inverse of $f$.
Since right- and left-moving conformal transformations commute, one may restrict attention to a single chirality, i.e. deal with $f$ alone (rather than a pair $f,\bar{f}$). Time translations and rotations then take the same form $f(\varphi)=\varphi+\text{const}$, as follows e.g. from the definition of the right-moving light-cone coordinate $\varphi^{-}$ in Eq. (2.1). This applies both to chiral systems and to nonchiral systems with periodic boundary conditions. (For nonchiral systems with e.g. Neumann boundary conditions, right and left movers are coupled at the boundaries, but the theory can be unfolded to a chiral one on a cylinder with twice the circumference.)
@@
${\cal U}(f)\,{\cal U}(g)=e^{ic\hskip 0.7pt\mathsf{C}(f,g)}\,{\cal U}(f\circ g),$ (2.3)
@@
$\mathsf{C}(f,g)\equiv\oint\frac{\textrm{d}\varphi}{48\pi}\,\log\bigl[f^{\prime}(g(\varphi))\bigr]\frac{g^{\prime\prime}(\varphi)}{g^{\prime}(\varphi)}$ (2.4)
@@
$\mathbf{X}=\sum_{m\in\mathbb{Z}}\hat{X}_{m}e^{im\varphi}\partial_{\varphi}.$ (2.5)
@@
$\mathfrak{u}(\mathbf{X})\equiv\partial_{\epsilon}\,{\cal U}(\mathbb{I}+\epsilon\mathbf{X})\Big|_{\epsilon=0}\equiv-i\oint\textrm{d}\varphi\,X(\varphi){\mathcal{T}}(\varphi),$ (2.6)
which defines the (Hermitian) stress-energy tensor ${\mathcal{T}}(\varphi)$. The latter is indeed the generator of conformal transformations, and may be seen as an operator-valued distribution ${\mathcal{T}}(\theta)=i\mathfrak{u}(\delta(\varphi-\theta)\partial_{\varphi})$, with $\delta(\varphi)$ the $2\pi$-periodic delta function. Expanded in Fourier modes,
${\mathcal{T}}(\varphi)=\frac{1}{2\pi}\sum_{n\in\mathbb{Z}}e^{in\varphi}\left(L_{n}-\frac{c}{24}\delta_{n,0}\right)\qquad\text{with}\qquad L_{m}\equiv i\mathfrak{u}\bigl(e^{-im\varphi}\partial_{\varphi}\bigr)+\frac{c}{24}\delta_{m,0},$ (2.7)
using which Eq. (2.6) becomes
$\mathfrak{u}(\mathbf{X})=-i\sum_{m}\hat{X}_{m}\left(L_{-m}-\frac{c}{24}\delta_{m,0}\right).$ (2.8)
The operators $L_{m}$ obtained in this way satisfy $L_{m}^{\dagger}=L_{-m}^{\vphantom{\dagger}}$ and the commutation relations
$L_{m},L_{n}]=(m-n)L_{m+n}+\frac{c}{12}m(m^{2}-1)\delta_{m+n,0}$ (2.9)
of the Virasoro algebra, consistent with the projective composition law (2.3) (see Appendix A).
2.2 Unitary drives in CFT
@@
In any (chiral) 1+1D CFT, time translations are generated by the zero mode of the stress-energy tensor (2.7), so the standard CFT Hamiltonian is
$H_{0}=\frac{2\pi}{L}\oint\textrm{d}\varphi\,{\mathcal{T}}(\varphi)=\frac{2\pi}{L}\Bigl(L_{0}-\frac{c}{24}\Bigr).$ (2.10)
@@
We refer to it as the undeformed or homogeneous Hamiltonian, in the sense that it is the Hamiltonian of the theory viewed from a suitable ‘rest frame’. Indeed, in CFT, observers are free to perform conformal changes of coordinates (2.1) and measure energy from a different ‘conformal frame’. (This is similar to special relativity, where Lorentz boosts affect the value of energy.) We therefore refer to the Hamiltonian in a different frame labeled by $f\in\operatorname{Diff}(S^{1})$, namely
$H[f]\equiv\,{\cal U}(f)H_{0}\,{\cal U}(f)^{-1},$ (2.11)
@@
as the CFT Hamiltonian deformed by $f$. This construction forms our starting point: viewing each operator (2.11) as being labeled by a conformal frame $f\in\operatorname{Diff}(S^{1})$, we will investigate the quantum geometry of the resulting bundle of eigenstates. The underlying parameter space is a quotient
${\cal M}\cong\operatorname{Diff}(S^{1})/S^{1},$ (2.12)
@@
since the undeformed Hamiltonian (2.10) commutes with $L_{0}$ but fails to commute with any other Virasoro generator, as is clear from the algebra (2.9). [^6] The notation $\operatorname{Diff}(S^{1})/S^{1}$ is standard for the space of all $S^{1}$ diffeomorphisms modulo rotations (see e.g. [57]). Technically, we should instead write $\operatorname{Diff}(S^{1})/\mathbb{R}$, since our notation $\operatorname{Diff}(S^{1})$’ refers to the universal cover of the diffeomorphism group of the circle. We stick to the standard notation nevertheless.
@@
As mentioned in Sec. 1, our strategy is to probe the geometry of parameter space by considering conformal transformations $f_{t}(\varphi)$ that vary in time $t$. The result is a CFT governed by a time-dependent deformed Hamiltonian
$H[f_{t}]=\,{\cal U}(f_{t})H_{0}\,{\cal U}(f_{t})^{-1}=\frac{2\pi}{L}\oint\textrm{d}\varphi\,\,{\cal U}(f_{t}){\mathcal{T}}(\varphi)\,{\cal U}(f_{t})^{-1},$ (2.13)
@@
where we used Eq. (2.10). This operator can be made explicit by relying on the usual transformation law of the CFT stress-energy tensor,
${\cal U}(f){\mathcal{T}}(\varphi)\,{\cal U}(f)^{-1}=[f^{\prime}(\varphi)]^{2}{\mathcal{T}}(f(\varphi))-\frac{c}{24\pi}\bigl\{f,\varphi\bigr\}$ (2.14)
@@
involving the Schwarzian derivative $\bigl\{f,\varphi\bigr\}\equiv\frac{f^{\prime\prime\prime}}{f^{\prime}}-\frac{3}{2}\Bigl(\frac{f^{\prime\prime}}{f^{\prime}}\Bigr)^{2}$. Plugging this into Eq. (2.13) yields
$H[f_{t}]=\frac{2\pi}{L}\oint\frac{\textrm{d}\varphi}{(f_{t}^{-1})^{\prime}(\varphi)}{\mathcal{T}}(\varphi)+C[f_{t}],$ (2.15)
@@
where $C[f]\equiv-\tfrac{c}{12L}\oint\textrm{d}\varphi\,\{f,\varphi\}$ is a real number. The key point is that Eq. (2.15) takes the form of a time-dependent inhomogeneous CFT (1.1), with the (dimensionless) velocity profile $v_{t}(x)\equiv V_{t}(\varphi)=1/(f_{t}^{-1})^{\prime}(\varphi)$ and the identification ${\mathcal{T}}(x)\equiv(2\pi/L)^{2}{\mathcal{T}}(\varphi)$ for $\varphi=2\pi x/L$. Clearly, such deformations of the homogeneous theory (2.10) leave a huge amount of legroom: by suitably choosing the maps $f_{t}(\varphi)$, one can obtain essentially any profile $v_{t}(x)$. The only requirements are $v_{t}(x)>0$ and the normalization $\frac{1}{L}\int_{0}^{L}\frac{\textrm{d}x}{v_{t}(x)}=\frac{1}{2\pi}\oint\frac{\textrm{d}\varphi}{V_{t}(\varphi)}=1$, due to the conditions (2.2).
@@
$L_{0}|h\rangle=h|h\rangle\qquad\text{and}\qquad L_{m}|h\rangle=0\quad\text{for all}\;\,m>0.$ (2.16)
Such a state can be obtained from the CFT vacuum $|0\rangle$, defined by $L_{m}|0\rangle=0$ for all $m\geq-1$, by acting on it with the corresponding primary field; see e.g. [58]. For our purposes, we may assume that there is at least one primary state in the theory, as our results can equally well be stated for the vacuum by setting $h=0$.
Given a primary state $|h\rangle$, it is standard practice to build a highest-weight representation of the Virasoro algebra (2.9), analogously to highest-weight representations of simple Lie algebras. [^7] Technically, the representation is lowest-weight, but the highest-weight terminology is standard. We assume that the representation is unitary, so that $h\geq 0$. Within that representation, the spectrum of $L_{0}$ is $\{h,h+1,h+2,h+3,\ldots\}$, so $|h\rangle$ is effectively a nondegenerate ground state with a (finite-volume) gap separating it from higher-energy states. Our bundle of Virasoro coherent states can then be constructed by acting on $|h\rangle$ with finite conformal transformations:
${\cal U}(f)|h\rangle\quad\text{for}\;\,f\in\operatorname{Diff}(S^{1}).$ (2.17)
Each of these is a normalized, nondegenerate eigenstate of the deformed Hamiltonian (2.11) with energy $(2\pi/L)(h-c/24)$ and a (finite-volume) gap above it. Such states are coherent in the same sense as Bloch coherent states or squeezed states in quantum optics [46], where the group $\operatorname{Diff}(S^{1})$ would respectively be replaced by $\mathrm{SU}(2)$ or $\mathrm{SL}(2,\mathbb{R})$.
2.3 Quantum geometric tensor in CFT
Having introduced the family of coherent states (2.17), let us write their quantum geometric tensor. The latter consists of a metric and a Berry curvature. We first recall the definition of these concepts in a broader setting, then specialize to the Virasoro case.
@@ 2.3 Quantum geometric tensor in CFT
${\cal M}\to\mathbb{P}{\cal H}:p\mapsto[|\phi(p)\rangle],$ (2.18)
@@
${\cal G}=\langle\textrm{d}\phi|\odot|\textrm{d}\phi\rangle-\langle\textrm{d}\phi|\phi\rangle\odot\langle\phi|\textrm{d}\phi\rangle,\qquad{\cal F}=-2i\langle\textrm{d}\phi|\wedge|\textrm{d}\phi\rangle,$ (2.19)
where $T_{1}\odot T_{2}\equiv(T_{1}\otimes T_{2}+T_{2}\otimes T_{1})/2$ and $T_{1}\wedge T_{2}\equiv(T_{1}\otimes T_{2}-T_{2}\otimes T_{1})/2$ respectively denote the symmetrized and antisymmetrized tensor products.
@@
Now choose some (isolated, nondegenerate) eigenstate $|h\rangle$ of $H_{0}$, and replace the parameter-dependent eigenstates $|\phi(p)\rangle$ in (2.18) by the states ${\cal U}(f)|h\rangle$, as in Eq. (2.17). The resulting metric and Berry curvature (2.19) are readily found thanks to the unitarity relation ${\cal U}(f)^{-1}=\,{\cal U}(f)^{\dagger}$, and they can be expressed in terms of the Lie algebra representation defined similarly to Eq. (2.6), namely $\mathfrak{u}(\mathbf{X})\equiv\partial_{\epsilon}\,{\cal U}(\mathbb{I}+\epsilon\mathbf{X})|_{\epsilon=0}$ for $\mathbf{X}\in\mathfrak{g}$. When evaluated at the identity coset labeled by $f=\mathbb{I}$ in parameter space (2.20), the metric and the curvature with arguments $\mathbf{X},\mathbf{Y}\in\mathfrak{g}$ read
${\cal G}(\mathbf{X},\mathbf{Y})$ $=-\frac{1}{2}\langle h|\bigl\{\mathfrak{u}(\mathbf{X}),\mathfrak{u}(\mathbf{Y})\bigr\}|h\rangle+\langle h|\mathfrak{u}(\mathbf{X})|h\rangle\langle h|\mathfrak{u}(\mathbf{Y})|h\rangle,$ (2.21a)
${\cal F}(\mathbf{X},\mathbf{Y})$ $=i\langle h|\bigl[\mathfrak{u}(\mathbf{X}),\mathfrak{u}(\mathbf{Y})\bigr]|h\rangle,$ (2.21b)
@@
The quantum metric and the Berry curvature can similarly be evaluated at any other point $fG_{0}$ in parameter space (2.20). In fact, they are fully determined by their expression (2.21) at the identity coset. The only difference is that the Lie algebra elements $\mathbf{X},\mathbf{Y}$ of Eq. (2.21) are then replaced by images of the Maurer-Cartan form acting on tangent vectors at $f$. This is to say that the quantum geometric tensor is invariant under $G$, which here stems from the form ${\cal U}(f)H_{0}\,{\cal U}(f)^{-1}$ of the deformed Hamiltonian, with unitary $\cal U$}. See Appendix B for details, which we omit here since these expressions will seldom be needed below.
Virasoro quantum geometry.
The formulas in Eq. (2.21) readily apply to the Virasoro coherent states of Sec. 2.2. In practice, the only complication is that the group $G=\operatorname{Diff}(S^{1})$ is now infinite-dimensional, so any tensor field on $G$ is a functional of a function $f(\varphi)$. What simplifies matters is that these tensor fields are invariant under $\operatorname{Diff}(S^{1})$, hence wholly determined by their expression at the identity $\mathbb{I}(\varphi)=\varphi$. For this reason, we now describe the quantum geometric tensor of CFTs by focusing entirely on Eq. (2.21) at the identity. Minimal effort is then needed to extend the result to fully fledged tensor fields on the entire group manifold, but this last bit is less essential for us and is relegated to Appendix B for brevity.
Consider first the ingredients needed in Eq. (2.21): $\mathbf{X},\mathbf{Y}$ are vector fields on the circle, the Lie algebra representation $\mathfrak{u}$ is the one in Eq. (2.6), and $|h\rangle$ is the highest-weight state introduced in Eq. (2.16). Fourier expanding $\mathfrak{u}(\mathbf{X}),\mathfrak{u}(\mathbf{Y})$ as in Eq. (2.7) and using the highest-weight conditions (2.16) then yields
${\cal G}(\mathbf{X},\mathbf{Y})$ $=\sum_{m\in\mathbb{Z}}|m|\Big(h+\frac{c}{24}(m^{2}-1)\Big)\hat{X}_{-m}\hat{Y}_{m}\equiv\sum_{m,n\in\mathbb{Z}}{\cal G}_{m,n}\hat{X}_{m}\hat{Y}_{n},$ (2.22a)
${\cal F}(\mathbf{X},\mathbf{Y})$ $=-2i\sum_{m\in\mathbb{Z}}m\Big(h+\frac{c}{24}(m^{2}-1)\Big)\hat{X}_{-m}\hat{Y}_{m}\equiv\sum_{m,n\in\mathbb{Z}}{\cal F}_{m,n}\hat{X}_{m}\hat{Y}_{n},$ (2.22b)

### v2

@@ I.3 Plan of the paper
The organization of the paper reflects the thematic split outlined above. In Sec. II, we introduce the family of CFT Hamiltonians (1) and their quantum geometry. Sec. III relates this to observable results, such as transition rates and linear-response coefficients, for small perturbations (3). Sec. IV concerns large but adiabatically driven deformations, and the general behavior of the return probability exemplified in Fig. 1 is derived for the important class of rotating drives. In Sec. V, we present families of time-dependent deformations that can be solved exactly using finite-dimensional matrices, without resorting to expansions in different limiting regimes. We conclude in Sec. VI, listing some natural extensions and follow-ups. Some technical details are deferred to appendices.
II Virasoro quantum geometry
@@
This section introduces the quantum geometry of the emergent conformal symmetry of gapless quantum systems, phrased within the context of CFT. We first recall some basics of conformal transformations and Virasoro group theory, then introduce inhomogeneous CFT and coherent states, and finally turn to our main theme: the quantum geometry of CFT.
II.1 Virasoro group theory
@@
We begin by reviewing conformal transformations in any 1+1D CFT. The presentation starts from the group of conformal transformations and introduces their algebra as a second step. [^5] Due to its infinite dimension, representations of the Virasoro group should strictly speaking be built the other way around: starting from the algebra and integrating to the group [78, 79]. We deliberately deviate from this, since our presentation generalizes directly to any quantum system with a (finite-dimensional) continuous symmetry group. Select formulas are summarized in Appendix A.
@@
$\varphi^{-}\mapsto f(\varphi^{-}),\qquad\varphi^{+}\mapsto\bar{f}(\varphi^{+}),$ (7)
@@
$f^{\prime}(\varphi)>0\qquad\text{and}\qquad f(\varphi+2\pi)=f(\varphi)+2\pi,$ (8)
and similarly for $\bar{f}$. The group operation in $\operatorname{Diff}(S^{1})$ is composition $(f\circ g)(\varphi)=f(g(\varphi))$, the identity element is $\mathbb{I}(\varphi)=\varphi$, and $f^{-1}$ denotes the inverse of $f$.
Since right- and left-moving conformal transformations commute, one may restrict attention to a single chirality, i.e. deal with $f$ alone (rather than a pair $f,\bar{f}$). Time translations and rotations then take the same form $f(\varphi)=\varphi+\text{const}$, as follows e.g. from the definition of the right-moving light-cone coordinate $\varphi^{-}$ in Eq. (7). This applies both to chiral systems and to nonchiral systems with periodic boundary conditions. (For nonchiral systems with e.g. Neumann boundary conditions, right and left movers are coupled at the boundaries, but the theory can be unfolded to a chiral one on a cylinder with twice the circumference.) We briefly return to this discussion at the end of Sec. IV, to explain how right and left chiralities combine in a driven CFT.
@@
${\cal U}(f)\,{\cal U}(g)=e^{ic\hskip 0.7pt\mathsf{C}(f,g)}\,{\cal U}(f\circ g),$ (9)
@@
$\mathsf{C}(f,g)\equiv\oint\frac{\textrm{d}\varphi}{48\pi}\,\log\bigl[f^{\prime}(g(\varphi))\bigr]\frac{g^{\prime\prime}(\varphi)}{g^{\prime}(\varphi)}$ (10)
@@
$\mathbf{X}=\sum_{m\in\mathbb{Z}}\hat{X}_{m}e^{im\varphi}\partial_{\varphi}.$ (11)
@@
$\mathfrak{u}(\mathbf{X})\equiv\partial_{\epsilon}\,{\cal U}(\mathbb{I}+\epsilon\mathbf{X})\Big|_{\epsilon=0}\equiv-i\oint\textrm{d}\varphi\,X(\varphi){\mathcal{T}}(\varphi),$ (12)
where we introduced the Hermitian operator-valued distribution ${\mathcal{T}}(\theta)=i\mathfrak{u}(\delta(\varphi-\theta)\partial_{\varphi})$ in terms of the $2\pi$-periodic delta function $\delta(\varphi)$. We take this $\mathcal{T}$} to be (the right-moving component of) the stress-energy tensor, as the latter is the generator of conformal transformations. Expanded in Fourier modes,
${\mathcal{T}}(\varphi)=\frac{1}{2\pi}\sum_{n\in\mathbb{Z}}e^{in\varphi}\left(L_{n}-\frac{c}{24}\delta_{n,0}\right)$ (13)
with $L_{n}\equiv i\mathfrak{u}\bigl(e^{-in\varphi}\partial_{\varphi}\bigr)+\frac{c}{24}\delta_{n,0}$, in terms of which Eqs. (11)–(12) yield
$\mathfrak{u}(\mathbf{X})=-i\sum_{m}\hat{X}_{m}\left(L_{-m}-\frac{c}{24}\delta_{m,0}\right).$ (14)
The operators $L_{n}$ obtained in this way satisfy $L_{n}^{\dagger}=L_{-n}^{\vphantom{\dagger}}$ and the commutation relations
$L_{m},L_{n}]=(m-n)L_{m+n}+\frac{c}{12}m(m^{2}-1)\delta_{m+n,0}$ (15)
of the Virasoro algebra, consistent with the projective composition law (9); see Appendix A for details.
II.2 Unitary drives in CFT
@@
In any (chiral) 1+1D CFT, time translations are generated by the zero mode of the stress-energy tensor, so the standard CFT Hamiltonian is
$H_{0}=\frac{2\pi}{L}\oint\textrm{d}\varphi\,{\mathcal{T}}(\varphi)=\frac{2\pi}{L}\Bigl(L_{0}-\frac{c}{24}\Bigr).$ (16)
@@
We refer to it as the undeformed or homogeneous Hamiltonian, in the sense that it is the Hamiltonian of the theory viewed from a suitable ‘rest frame’. Indeed, in CFT, observers are free to perform conformal changes of coordinates (7) and measure energy from a different ‘conformal frame’. (This is similar to special relativity, where Lorentz boosts affect the value of energy.) We therefore refer to the Hamiltonian in a different frame labeled by $f\in\operatorname{Diff}(S^{1})$, namely
$H[f]\equiv\,{\cal U}(f)H_{0}\,{\cal U}(f)^{-1},$ (17)
@@
as the CFT Hamiltonian deformed by $f$. This construction forms our starting point: viewing each operator (17) as being labeled by a conformal frame $f\in\operatorname{Diff}(S^{1})$, we will investigate the quantum geometry of the resulting bundle of eigenstates. The underlying parameter space is a quotient
${\cal M}\cong\operatorname{Diff}(S^{1})/S^{1},$ (18)
@@
since the undeformed Hamiltonian (16) commutes with $L_{0}$ but fails to commute with any other Virasoro generator, as is clear from the algebra (15). [^7] The notation $\operatorname{Diff}(S^{1})/S^{1}$ is standard for the space of all $S^{1}$ diffeomorphisms modulo rotations (see e.g. [57]). Technically, we should instead write $\operatorname{Diff}(S^{1})/\mathbb{R}$, since our notation $\operatorname{Diff}(S^{1})$’ refers to the universal cover of the diffeomorphism group of the circle. Topologically, the parameter space (18) is homeomorphic to an (infinite-dimensional) linear space, so it is homotopic to a point; see e.g. [81, 82].
@@
As mentioned in Sec. I, our strategy is to probe the geometry of parameter space by considering conformal transformations $f_{t}\in\operatorname{Diff}(S^{1})$ that vary continuously in time $t$. The result is a CFT governed by a time-dependent deformed Hamiltonian
$H[f_{t}]$ $=\,{\cal U}(f_{t})H_{0}\,{\cal U}(f_{t})^{-1}$ (19)
$=\frac{2\pi}{L}\oint\textrm{d}\varphi\,\,{\cal U}(f_{t}){\mathcal{T}}(\varphi)\,{\cal U}(f_{t})^{-1},$
@@
where we used Eq. (16). This operator can be made explicit by relying on the usual transformation law of the CFT stress-energy tensor,
${\cal U}(f){\mathcal{T}}(\varphi)\,{\cal U}(f)^{-1}=[f^{\prime}(\varphi)]^{2}{\mathcal{T}}(f(\varphi))-\frac{c}{24\pi}\bigl\{f,\varphi\bigr\},$ (20)
@@
with $\bigl\{f,\varphi\bigr\}\equiv{f^{\prime\prime\prime}}/{f^{\prime}}-\frac{3}{2}\bigl({f^{\prime\prime}}/{f^{\prime}}\bigr)^{2}$ the Schwarzian derivative of $f$. Plugging this into Eq. (19) yields
$H[f_{t}]=\frac{2\pi}{L}\oint\frac{\textrm{d}\varphi}{(f_{t}^{-1})^{\prime}(\varphi)}{\mathcal{T}}(\varphi)+C[f_{t}],$ (21)
@@
where $C[f]\equiv-\tfrac{c}{12L}\oint\textrm{d}\varphi\,\{f,\varphi\}$ is a real number. The key point is that Eq. (21) takes the form of a time-dependent inhomogeneous CFT (1), with the (dimensionless) velocity $v_{t}(x)\equiv V_{t}(\varphi)=1/(f_{t}^{-1})^{\prime}(\varphi)$ and the identification ${\mathcal{T}}(x)\equiv(2\pi/L)^{2}{\mathcal{T}}(\varphi)$ for $\varphi=2\pi x/L$. Clearly, such deformations of the homogeneous theory (16) leave a huge amount of legroom: by suitably choosing the maps $f_{t}(\varphi)$, one can obtain essentially any profile $v_{t}(x)$. The only requirements are $v_{t}(x)>0$ and the normalization $\frac{1}{L}\int_{0}^{L}\frac{\textrm{d}x}{v_{t}(x)}=\frac{1}{2\pi}\oint\frac{\textrm{d}\varphi}{V_{t}(\varphi)}=1$, due to the conditions (8).
@@
$L_{0}|h\rangle=h|h\rangle,\quad L_{m}|h\rangle=0\quad\text{for all}\;\,m>0.$ (22)
Such a state can be obtained from the CFT vacuum $|0\rangle$, which satisfies $L_{m}|0\rangle=0$ for all $m\geq-1$, by acting on it with the corresponding primary field; see e.g. [83]. For our purposes, we may assume that there is at least one primary state in the theory, as our results can equally well be stated for the vacuum by setting $h=0$. This is important for our numerical simulations, all of which involve the ground state $|h\rangle=|0\rangle$.
Given a primary state $|h\rangle$, it is standard practice to build a highest-weight representation of the Virasoro algebra (15), analogously to highest-weight representations of simple Lie algebras. [^8] Technically, the representation is lowest-weight, but the highest-weight terminology is standard. We assume that the representation is unitary, so that $h\geq 0$. Within that representation, the spectrum of $L_{0}$ is (a subset of) $\{h,h+1,h+2,h+3,\ldots\}$, so $|h\rangle$ is effectively a nondegenerate ground state with a (finite-volume) gap separating it from higher-energy states. Our bundle of Virasoro coherent states can then be constructed by acting on $|h\rangle$ with finite conformal transformations:
${\cal U}(f)|h\rangle\quad\text{for}\;\,f\in\operatorname{Diff}(S^{1}).$ (23)
Each of these is a normalized, nondegenerate eigenstate of the deformed Hamiltonian (17) with energy $(2\pi/L)(h-c/24)$ and a (finite-volume) gap above it. Such states are coherent in the same sense as Bloch coherent states or squeezed states in quantum optics [84, 67], where the group $\operatorname{Diff}(S^{1})$ would respectively be replaced by $\mathrm{SU}(2)$ or $\mathrm{SL}(2,\mathbb{R})$.
II.3 Quantum geometric tensor
Having introduced the family of coherent states (23), let us write their quantum geometric tensor. The latter consists of a metric and a Berry curvature. We first recall the definition of these concepts in a broader setting, then specialize to the Virasoro case.
@@ II.3 Quantum geometric tensor
${\cal M}\to\mathbb{P}{\cal H}:p\mapsto\bigl[|\phi(p)\rangle\bigr],$ (24)
@@
${\cal G}$ $=\langle\textrm{d}\phi|\odot|\textrm{d}\phi\rangle-\langle\textrm{d}\phi|\phi\rangle\odot\langle\phi|\textrm{d}\phi\rangle,$ (25a)
${\cal F}$ $=-2i\langle\textrm{d}\phi|\wedge|\textrm{d}\phi\rangle,$ (25b)
where $T_{1}\odot T_{2}\equiv(T_{1}\otimes T_{2}+T_{2}\otimes T_{1})/2$ and $T_{1}\wedge T_{2}\equiv(T_{1}\otimes T_{2}-T_{2}\otimes T_{1})/2$ respectively denote the symmetrized and antisymmetrized tensor products. The pair $({\cal F},{\cal G})$ is compatible in the sense of Kähler manifolds when the map (24) is not only smooth but analytic [85], which will turn out to be the case for the Virasoro coherent states (23).
@@
Now choose some (isolated, nondegenerate) eigenstate $|h\rangle$ of $H_{0}$, and replace the parameter-dependent eigenstates $|\phi(p)\rangle$ in (24) by the states ${\cal U}(f)|h\rangle$, as in Eq. (23). The resulting metric and Berry curvature (25) are readily found thanks to the unitarity relation ${\cal U}(f)^{-1}=\,{\cal U}(f)^{\dagger}$, and they can be expressed in terms of the Lie algebra representation defined similarly to Eq. (12), namely $\mathfrak{u}(\mathbf{X})\equiv\partial_{\epsilon}\,{\cal U}(\mathbb{I}+\epsilon\mathbf{X})|_{\epsilon=0}$ for $\mathbf{X}\in\mathfrak{g}$. When evaluated at the identity coset labeled by $f=\mathbb{I}$ in parameter space (26), the metric and the curvature with arguments $\mathbf{X},\mathbf{Y}\in\mathfrak{g}$ read
${\cal G}(\mathbf{X},\mathbf{Y})$ $=-\frac{1}{2}\langle h|\bigl\{\mathfrak{u}(\mathbf{X}),\mathfrak{u}(\mathbf{Y})\bigr\}|h\rangle$
$\quad\,+\langle h|\mathfrak{u}(\mathbf{X})|h\rangle\langle h|\mathfrak{u}(\mathbf{Y})|h\rangle,$ (27a)
${\cal F}(\mathbf{X},\mathbf{Y})$ $=i\langle h|\bigl[\mathfrak{u}(\mathbf{X}),\mathfrak{u}(\mathbf{Y})\bigr]|h\rangle,$ (27b)
@@
${\cal G}(\mathbf{X},\mathbf{X})=-\langle h|\mathfrak{u}(\mathbf{X})^{2}|h\rangle+\langle h|\mathfrak{u}(\mathbf{X})|h\rangle^{2},$ (28)
which is nothing but the variance of the Hermitian operator $i\mathfrak{u}(\mathbf{X})$. Note that both the metric and the curvature are generally degenerate on the manifold (26), though they can be nondegenerate on further quotient spaces thereof: we will see examples of this below. (No such complications arise for Bloch coherent states or squeezed states.)
The quantum metric and the Berry curvature can similarly be evaluated at any other point $fG_{0}$ in parameter space (26). In fact, they are fully determined by their expressions (27) at the identity. The only difference is that the Lie algebra elements $\mathbf{X}$ and $\mathbf{Y}$ are then replaced by images of the Maurer-Cartan form acting on tangent vectors at $f$. This is to say that the quantum geometric tensor is invariant under $G$, which stems here from the form ${\cal U}(f)H_{0}\,{\cal U}(f)^{-1}$ of the deformed Hamiltonian. See Appendix B for further details, which we omit here since they will seldom be needed below.
Virasoro case.
[…]
Consider first the ingredients needed in Eq. (27): $\mathbf{X},\mathbf{Y}$ are vector fields on the circle, the Lie algebra representation $\mathfrak{u}$ is the one in Eq. (12), and $|h\rangle$ is the highest-weight state introduced in Eq. (22). Fourier expanding $\mathfrak{u}(\mathbf{X}),\mathfrak{u}(\mathbf{Y})$ as in Eq. (13) and using the highest-weight conditions (22) then yields
${\cal G}(\mathbf{X},\mathbf{Y})$ $=\sum_{m\in\mathbb{Z}}|m|\Big(h+\frac{c}{24}(m^{2}-1)\Big)\hat{X}_{-m}\hat{Y}_{m}$
$\equiv\sum_{m,n\in\mathbb{Z}}{\cal G}_{m,n}\hat{X}_{m}\hat{Y}_{n},$ (29a)
${\cal F}(\mathbf{X},\mathbf{Y})$ $=-2i\sum_{m\in\mathbb{Z}}m\Big(h+\frac{c}{24}(m^{2}-1)\Big)\hat{X}_{-m}\hat{Y}_{m}$
$\equiv\sum_{m,n\in\mathbb{Z}}{\cal F}_{m,n}\hat{X}_{m}\hat{Y}_{n},$ (29b)